Suppose eal GDP is growing at 4 percent, the money supply is growing at 11 percent, the velocity of money is constant, and the real interest is 6 percent.
A) what is the current inflation rate and nominal interest rate?
B) if the money supply growth rate increases to 15 percent, how will your answer in part (a) change?
C) if you were an investor, how would the change in the money supply growth affect your real profiability, assuming that you now receive the new nominal interest rate?
D) based on your prwvious answers, would you prefer a fixed or a floating interest on your investment?
1
Expert's answer
2012-09-07T09:35:59-0400
i) M/P=Y/V
Y - growing by 4% so coeficient = 1.04
M - growing by 11% so coeficient = 1.11
V - constant =1
1.11/P=1.04/1
P=1.11/1.04≈1.067≈+6.7%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.067)
i=(1.06*1.067)-1
i≈0.13102≈13.10%
______ ______ ______ ______

(ii)
M/P=Y/V
Y - growing by 4% so coeficient = 1.04
M - growing by 15% so coeficient = 1.15
V - constant =1
1.15/P=1.04/1
P=1.15/1.04≈1.1057692≈+10.57%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.1057)
i=(1.06*1.1057)-1
i≈0.1720≈17.20%
______ ______ ______ ______

i¹≈13.10%
i²≈17.20%
Δi=(17.20/13.10)-1≈
≈1.3129-1≈+31.29%